suppose that the demand function or schedule is given by the linear equation p = 300 - qd and the supply equation by p =60 + 2 qs (a) find the equilibrium price and quantity (b) suppose the supply curve slop prices from 2 qs to 3 qs (C) find the new equilibrium price and quantity (d) find the equilibrium position graphically?
1
Expert's answer
2021-07-30T02:53:02-0400

"Qd=300-p\\
Qs=p-60\div2"

in equation

"Qd=Qs"

"(300-p)=p-60\div2"

"600-2p=p-60\\
600+60=p+2p"

"660=3p\\
p=110"


(b)

"p=60+3Qs\\
Qs=p-60\div3\\
Qd=300-p\\
Qs=Qd\\
p-60\div3=300-p\\
p-60=900-3p\\
4p=900+60\\
p=960\div4\\
p=240"